OpenStudy (anonymous):

a simple gear train is composed of three gears. gear a is the driver and has 22 teeth gear b has 15 teeth and gear c has 84 if the output is at C what is the gear ratio?

OpenStudy (anonymous):

N- Speed of gears, Z- No. of gear teeth Gear ratio = Speed of input gear to speed of output gear Na/Nc = (Nb/Nc)*(Na/Nb) = (Zc/Zb)*(Zb/Za) = (84/15)*(15/22) = 4 Gear ratio = 4